# Hybrid Video Diffusion Models with 2D Triplane and 3D Wavelet Representation

### v1

@@ 4.2 Loss function
Figure 4: Qualitative reconstruction results on UCF-101 [35], SkyTimelapse [46], and TaiChi dataset [30]. Our HVDM faithfully express a complex structures with fine details. Additional samples are included in appendix.
4.3 Diffusion-based Video Generator
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ and $\beta$ represents a predefined schedule for the variance.
where $\epsilon_{\theta}$ is a trainable denoising diffusion model which recover a noisy sample $\mathbf{z}_{t}$. The denoising diffusion model is trained to predict the noise with the following simplified objective:
4.4 Applications for Diverse Video Generator
VideoGPT MoCoGAN + StyleGAN2 MoCoGAN-HD DIGAN StyleGAN-V ${}^{\dagger}$ ${}^{\dagger}$ HVDM (ours)
@@ 4.4 Applications for Diverse Video Generator
${}^{\dagger}$ ${}^{\dagger}$ HVDM (ours)
@@
${}^{\dagger}$ $*{}^{*}$ HVDM (ours)
@@
Table 2: Quantitative comparison of video generation models on UCF-101, SkyTimelapse and TaiChi datasets. We evaluate FVD and KVD metrics (the lower values are better) and the best score are marked as bold. The overall score of other baselines are referenced from StyleGAN-V [33]. Methods marked with ${}^{\dagger}$ denote the score is computed using official pretrained weights and $*{}^{*}$ denote the score is reproduced using official codes. We report evaluation details and sampling settings on Appendix B.
Figure 5: Generated video sample from our HVDM on UCF-101 [35], SkyTimelapse [46], and TaiChi dataset [30].
Long video generation. To facilitate long video generation, our work follows the traditional frame prediction approach [41, 13, 48] where previous frame conditions are combined with future frames. Given two consecutive video clips $[\mathbf{x}^{\mathrm{prev}},\mathbf{x}^{\mathrm{next}}]$ of $\mathbf{x}^{\mathrm{prev}},\mathbf{x}^{\mathrm{next}}\in\mathbb{R}^{3\times S\times H\times W}$, we can train the model to generate initial clips and future clips in a unified manner. More specifically, we can generate an initial clip with a null latent and extend the video by composing the condition with previous frame latent as follow:
$\mathbf{z}_{c}^{\mathrm{long}}=\begin{cases}\mathcal{E}(\mathbf{x}^{\mathrm{prev}})&~{}\text{if future clip}\\ 0&~{}\text{otherwise.}\end{cases}$ (14)
where $\mathcal{E}$ denotes our pre-trained hybrid encoder.
Image-to-Video. Image-to-video generation aims to synthesize a coherent video starting from an image. In general, videos are composed of a series of consecutive images, where the first frame serves as a representative depiction of the overall content within the video. To facilitate the process of image-to-video generation without introducing an additional image encoder, we repeatedly stack the first frame of the video on the temporal axis. Then, we use the latent $\mathcal{I}^{\mathrm{content}}$ extracted from our video encoder as the content condition. By leveraging the latent $\mathcal{I}^{\mathrm{content}}$ which capture the common content, we can effectively guide the model to generate coherent video sequence as follow:
Video dynamics control. Empowering users to control motion intensity in video creation provides flexibility with an improved user experience. Unfortunately, it was difficult to obtain motion-related labels from existing video collections for training diffusion models. To enable the alteration of motion intensity, we devise a method that learns motion dynamics with inter-frame variance. Specifically, we define the level of motion alteration as the structural dissimilarity between the semantic latent of the start and end frames. Since the structural dissimilarity encompasses both semantic and motion variance, we employ it as guiding cues to control motion without explicit motion labeled video. Formally, our structural dissimilarity is defined as follow:
$\mathbf{z}_{c}^{\mathrm{dynamics}}=\|\mathbf{z}_{0}^{\tt T}-\mathbf{z}_{0}^{\tt 1}\|_{1}$ (16)
$\begin{split}\vspace{-0.02in}\mathbb{E}_{\mathbf{x}_{0},\epsilon,t}\Big{[}&\lambda||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{z}_{c},t)||_{2}^{2}\\ &+(1-\lambda)||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},{0},t)||_{2}^{2}\Big{]}.\end{split}\vspace{-0.02in}$ (17)
5 Experiments
5.1 Training Details
Datasets. We conduct the experiments on three commonly-used benchmark datasets for the video generation task: UCF-101 [35], SkyTimelapse [46], and TaiChi [30]. UCF-101 [35] contains 13,320 videos from 101 action classes, offering a diverse range of human actions. SkyTimelapse [46] is composed of 5,000 videos containing dynamic sky scenes. TaiChi [30] consists of full human bodies performing Tai-Chi actions.
Training. Our implementation is mainly based on PVDM [48]. The video autoencoder component for 2D projected latent stem from the PVDM [48], but the other wavelet-based frequency component is added. We train our models on 4 NVIDIA A100 GPUs for each autoencoding and diffusion process. See Appendix A for the details.
5.2 Evaluation Details
Baseline models. We compare our HVDM with the following video generation models: VideoGPT [47], MoCoGAN [37], MoCoGAN-HD [36], DIGAN  [49], StyleGAN-V [32], LVDM [13], and PVDM  [48]. We utilize official model weights and conduct training using the official code when the model weights are not available.
[…]
@@ 5.3 Comparison Results
Reconstruction results. Fig. 4 and Appendix D.1 show the performance of the reconstructed videos from our autoencoder and other diffusion-based video generative models. Our autoencoder presents a more accurate and clear reconstructed video compared to other baseline methods. In the Tab. 1, we demonstrate that our autoencoder achieves superior performance in all metrics for the three datasets.
@@
Generation results. We perform video generation on three datasets and compare the results with the other video generation methods. As shown in Fig. 5, our method is capable of generating not only monotonic motion videos such as SkyTimelapse [46] and TaiChi [30] but also dynamic motion video UCF-101 [35], which contain complex action categories. Note that synthesizing high quality videos on a complex dataset like UCF-101 [35] has been a challenging problem where existing video generation method often struggle with producing natural motions with good quality. As shown in Tab. 2, our method achieves superior performance in the benchmark datasets. See Appendix D for more samples.
Figure 6: Visualization of our hybrid representations. The original video frames (1st row) and the visualization of their corresponding projected latent (2nd row), low-frequency (3rd row), high-frequency (4th row), and hybrid latent (5th row) components are shown.
5.4 Analysis and Ablation Study
@@
Interpreting the frequency supervision. We perform visualization to understand how hybrid representation can help autoencoder to encode the 3D video pixels in Fig. 6. The decomposed 3D wavelet features is divided into low frequency and high frequency components, where each component conduct cross-attention operation with the 2D projected latent. As shown in Fig. 6, low frequency component captures static and dynamic regions and high frequency component capture the sharp edges and important details of video. To refine the ambiguity of the 2D project latent which contain global and abstract information, we incorporate these latent with frequency components. As a result, our autoencoder figures out where the objects are located and how object’s motion occurs in the video.
Ablation study. Ablation studies of our autoencoder on UCF-101 can be found in Appendix C. We conduct verification of performance based on the autoencoder’s backbone architecture in Appendix C.1. As shown in Appendix D.1, our hybrid autoencoder shows substantially better naturalness while maintaining the perceptual details of the video. In the Appendix C.2, we verify several techniques (e.g., concatenation or cross-attention) for integrating the frequency component with a 2D projected latent.
@@ 6 Conclusion
In this paper, we proposed a hybrid video autoencoder for video generation using diffusion models, HVDM. We investigated the effectiveness of our hybrid architecture that combines 2D projected and 3D volume representation, leveraging the strengths of frequency matching loss. By incorporating these representations with spatio-temporal cross-attention, HVDM generates high-quality videos with improved realism. Our extensive experiments on benchmark datasets demonstrated the superior performance, compared to the other baseline methods.
Appendix
Appendix A Implementation Details
In this section we describe details on the model configurations and hyper-parameters used for training the models. For 2D triplane and 3D wavelet representation of our hybrid autoencoder, we modify existing PVDM autoencoder architecture: 1) we use reduced depth of each layer in the transformer, 2) we use the decomposed wavelet input instead of the raw video, 3) we used 3D convolution networks to encode the decomposed wavelet features, and 4) we add cross-attention module to fuse 2D projected and 3D wavelet latent. As for the diffusion model, we use a simple 3D convolutional U-Net architecture. We provide the detailed information for our hybrid autoencoder in Tab. 3 and U-Net in Tab. 4.
UCF-101 SkyTimelapse TaiChi
@@ Appendix A Implementation Details
Table 3: Autoencoder configuration and training details. We combine transformer and 3D convolution networks for hybrid video representation.
UCF-101 SkyTimelapse TaiChi
@@
Table 4: Our denoising 3D U-Net configuration and training details. We train a conventional 3D U-Net as denoiser that uses the skip connections.
Appendix B Experiment Setup
B.1 Datasets
For fair comparison with baseline, we follow the settings of PVDM for UCF-101 and SkyTimelapse datasets and as well as the settings of LVDM for the TaiChi dataset. For training the model, we crop and resize all video clips to $256\times 256$ scale. All video clips are extracted as consecutive 16 or 128 frame lengths from a randomly selected video.
UCF-101 is an action recognition dataset comprising realistic action videos, including 101 action categories. Following the common practice for evaluating unconditional video generation on UCF-101, we use the train split (9,357 videos) for training and test split (3,963 videos) for evaluation.
@@ B.1 Datasets
TaiChi is a Tai-Chi action dataset containing 3,078 videos in total. For a fair comparison with the LVDM, we use the train+test splits for both model training and evaluation. Moreover, we reproduced PVDM [48] under the same setup as the model weight was unavailable.
B.2 Metrics
For fair comparison, we follows the evaluation protocol of StyleGAN-V and PVDM. Unlike previous evaluation protocol that preprocess video datasets with fixed-length clips, StyleGAN-V propose to sample video data first and then randomly select fixed-length clips, particularly avoiding bias issues with extremely long videos. As for evaluation metrics, we measure the video generation quality with the following two metrics: Fréchet Video Distance (FVD) and Kernel Video Distance (KVD). In addition, we report the reconstruction quality of the autoencoder with the following four metrics: Learned Perceptual Image Patch Similarity (LPIPS), Peak Signal-to-Noise Ratio (PSNR), SSIM (Structural Similarity Index), and R-FVD which denote FVD between real and reconstructed samples.
Inference settings for evaluations For the fair comparison between diffusion-based video generative models, we fixed the number of DDIM steps to 100. Additionally we extract 2,048 video clips for computing real statistics and generate 2,048 samples for evaluating fake statistics.
Appendix C Ablation Studies
C.1 Ablation study on backbone architectures
We perform an ablation study of backbone architectures on UCF-101 dataset. As shown in Tab. 5, our hybrid autoencoder demonstrates effectiveness in capturing both perceptual details and temporal coherence in the video. We can observe that 3D CNN-based autoencoder show better performance compared to the 2D triplane-based autoencoder since it has more latent dimensions. Although the 2D triplane-based autoencoder effectively compresses latent dimensions, enhancing temporal coherence necessitates adversarial training with a 3D discriminator. Our hybrid autoencoder integrates the strengths of both architectures, achieving superior performance.
Backbone Proj. Adv. $\mathbf{z}$ shape(dim) R-FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Video Transformer - $4\times(32\times 32+16\times 32+16\times 32)$ 180.21 0.092 28.12 0.782
Video Transformer $4\times(32\times 32+16\times 32+16\times 32)$ 027.03 0.095 27.03 0.742
3D CNN - - $4\times 16\times 32\times 32$ 086.73 0.057 31.76 0.852
3D CNN - $4\times 16\times 32\times 32$ 021.69 0.064 30.39 0.837
Ours - $4\times 16\times 32\times 32$ 005.35 0.038 34.00 0.915
Table 5: Ablation study on various backbone architectures
C.2 Ablation study on fusing modules
We conduct several experiments to assess how the fusing module affects the performance of our hybrid autoencoder. For feature fusion, we adopted two different modules: 1) the concatenation module 2) the cross-attention module. We also validated the effectiveness of wavelet decomposition by comparing the results which use raw RGB video as input. As shown in Tab. 7, the experimental results show that the cross-attention module and wavelet-decomposed components contribute to the improvement of reconstruction performance. We analyze that the cross-attention module combines different 2D triplane and 3D wavelet representations better and the wavelet-decomposed components are prompted for better spatio-temporal alignment by a frequency matching loss.
HVDM (w/o FML) 8.08 0.041 33.67 0.892
HVDM (w/0 FML) 5.35 0.038 34.00 0.915
Table 6: Ablation study on frequency matching loss.
C.3 Inference time and memory
@@
Fusion module R-FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ Baseline 27.03 0.095 27.03 0.742 Baseline + RGB (Concat.) 20.37 0.063 31.02 0.831 Baseline + Low Freq. (Concat.) 34.39 0.078 30.50 0.816 Baseline + High Freq. (Concat.) 14.96 0.061 31.32 0.838 Baseline + Low & High Freq. (Concat.) 09.92 0.051 32.69 0.871 Baseline + RGB (Cross-attn.) 06.32 0.038 33.79 0.908 Baseline + Low Freq. (Cross-attn.) 06.87 0.040 33.69 0.909 Baseline + High Freq. (Cross-attn.) 06.97 0.047 33.20 0.899 Ours 05.35 0.037 34.00 0.915 Table 7: Ablation study on feature fusing module.
Inference (time/memory) Length $\rightarrow$ 16 128 LVDM [13]; 8.85/.9.0 N/A PVDM [48]; 9.16/.9.6 137.8/10.1 HVDM (ours); 9.90/11.6 151.4/12.1 Table 8: Comparison of the inference time (s) and memory (GB) of various video diffusion models. N/A denotes that the values cannot be measured because the model is unavailable.
Appendix D ADDITIONAL RESULTS
D.1 VIDEO RECONSTRUCTION
Figure 7: Reconstruction results of our hybrid autoencoder on the SkyTimelapse.
Figure 8: Reconstruction results of our hybrid autoencoder on the TaiChi.
Figure 9: Reconstruction results of our hybrid autoencoder on the UCF-101.
D.2 UNCONDITIONAL VIDEO GENERATION

### v4

@@ 5.1 Training Details
Figure 4: Qualitative reconstruction results on SkyTimelapse [51] dataset [33]. Our HVDM produces distinct and sharp edge details and faithfully expresses a fine structures in a video sequence. Additional samples are included in appendix.
Training. The video autoencoder component for 2D projected latent stem from the PVDM [53], but the other wavelet-based frequency component is added. We train our models on 4 NVIDIA A100 GPUs for each autoencoding and diffusion process. See Appendix A for the details.
5.2 Evaluation Details
Baseline models. We compare HVDM with the following models: VideoGPT [52], MoCoGAN [40], MoCoGAN-HD [39], DIGAN [54], StyleGAN-V [35], LVDM [14], and PVDM  [53]. We utilize official model weights and conduct training using the official code when the model weights are not available.
Quantitative evaluations. To evaluate reconstruction quality, we use R-FVD [42], LPIPS [55], PSNR [21], and SSIM [19]. Here, R-FVD denotes FVD between reconstructed videos and the ground-truth videos. As for the evaluation of generation quality, we measure FVD [42] and KVD [41]. Considering different clip lengths, we denote FVD16 and KVD16 scores, as well as FVD128 and KVD128 scores for video clips with a length of 16 and 128. For a fair comparison, we follow the evaluation protocol of StyleGAN-V [36]; i.e., we use 2,048 real/fake video clips for evaluating FVD. See Appendix B for the details.
VideoGPT MoCoGAN + StyleGAN2 MoCoGAN-HD DIGAN StyleGAN-V LVDM PVDM HVDM (ours)
@@ 5.2 Evaluation Details
LVDM PVDM HVDM (ours)
@@
LVDM PVDM HVDM (ours)
@@
Table 2: Quantitative comparison of video generation models on UCF-101, SkyTimelapse and TaiChi datasets. We evaluate FVD and KVD metrics (the lower values are better) and the best score are marked as bold. The overall score of other baselines are referenced from StyleGAN-V [36]. Methods marked with denote the score is computed using official pretrained weights and denote the score is reproduced using official codes. We report evaluation details and sampling settings on Appendix B.
Figure 5: Generated video sample by our HVDM on UCF-101 [38], SkyTimelapse [51], and TaiChi datasets [33].
@@ 5.3 Comparison Results
Reconstruction results. Fig. 4 and Appendix E.1 show the performance of the reconstructed videos from our autoencoder and other diffusion-based video generative models. Our autoencoder presents a more accurate and clear reconstructed video compared to other baseline methods. In the Tab. 1, we demonstrate that our autoencoder achieves superior performance in all metrics for the three datasets.
@@
Generation results. We perform video generation on three datasets and compare the results with the other video generation methods. As shown in Fig. 5, our method is capable of generating not only monotonic motion videos such as SkyTimelapse [51] and TaiChi [33] but also dynamic motion video UCF-101 [38], which contain complex action categories. Note that synthesizing high quality videos on a complex dataset like UCF-101 [38] has been a challenging problem where existing video generation method often struggle with producing natural motions with good quality. As shown in Tab. 2, our method achieves superior performance in the benchmark datasets. See Appendix E for more samples.
5.4 Interpreting the frequency supervision
Figure 6: Visualization of our hybrid representations.
@@ 5.4 Interpreting the frequency supervision
We perform visualization to understand how hybrid representation can help autoencoder to encode the 3D video pixels in Fig. 6. The decomposed 3D wavelet features is divided into low frequency and high frequency components, where each component conduct cross-attention operation with the 2D projected latent. As shown in Fig. 6, low frequency component captures static and dynamic regions and high frequency component capture the sharp edges and important details of video. To refine the ambiguity of the 2D project latent which contain global and abstract information, we incorporate these latent with frequency components. As a result, our autoencoder figures out where the objects are located and how object’s motion occurs in the video.
5.5 Ablation study on backbone architectures
We conduct verification of performance based on the autoencoder’s backbone architecture on UCF-101 dataset. As shown in Tab. 3, our hybrid autoencoder demonstrates effectiveness in capturing both perceptual details and temporal coherence in the video. We can observe that 3D CNN-based autoencoder show better performance compared to the 2D triplane-based autoencoder since it has more latent dimensions. Although the 2D triplane-based autoencoder effectively compresses latent dimensions, enhancing temporal coherence necessitates adversarial training with a 3D discriminator. Our hybrid autoencoder integrates the strengths of both architectures, achieving superior performance.
Backbone Proj. Adv. $\mathbf{z}$ shape dimension R-FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Video Transformer [5] - $4\times(32\times 32+16\times 32+16\times 32)$ 180.21 0.092 28.12 0.782
Video Transformer [5] $4\times(32\times 32+16\times 32+16\times 32)$ 027.03 0.095 27.03 0.742
3D CNN [23] - - $4\times 16\times 32\times 32$ 086.73 0.057 31.76 0.852
3D CNN [23] - $4\times 16\times 32\times 32$ 021.69 0.064 30.39 0.837
Ours - $\bm{4\times 16\times 32\times 32}$ 005.35 0.038 34.00 0.915
Table 3: Ablation study on various backbone architectures
5.6 Ablation study on fusing modules
We verify several experiments to assess how the fusing module affects the performance of our hybrid autoencoder on UCF-101 dataset. For feature fusion, we adopted two different modules: 1) the concatenation module 2) the cross-attention module. We also validated the effectiveness of wavelet decomposition by comparing the results which use raw RGB video as input. As shown in Tab. 4, the experimental results show that the cross-attention module and wavelet-decomposed components contribute to the improvement of reconstruction performance. We analyze that utilizing the temporal and spatial attention module facilitate the integration of different 2D triplane and 3D wavelet representation through the comparison of features at different spatial locations.
Fusion module R-FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Baseline 27.03 0.095 27.03 0.742
Baseline + RGB (Concat.) 20.37 0.063 31.02 0.831
Baseline + Low Freq. (Concat.) 34.39 0.078 30.50 0.816
Baseline + High Freq. (Concat.) 14.96 0.061 31.32 0.838
Baseline + Low & High Freq. (Concat.) 09.92 0.051 32.69 0.871
Fusion module R-FVD $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
Baseline + RGB (Cross-attn.) 06.32 0.038 33.79 0.908
Baseline + Low Freq. (Cross-attn.) 06.87 0.040 33.69 0.909
Baseline + High Freq. (Cross-attn.) 06.97 0.047 33.20 0.899
Ours 05.35 0.037 34.00 0.915
Table 4: Ablation study on feature fusing module.
5.7 Ablation study on frequency matching loss
We validate the effectiveness of frequency matching loss (FML) on UCF-101 dataset in Tab. 5. The result shows that utilizing the FML term leads to a remarkable reduction in the FVD score by approximately 2.73 points demonstrating its effectiveness in enhancing the temporal coherency. Furthermore, evaluation metrics related to perceptual details also improved to 0.003 points for LPIPS, 0.33 points for PSNR, and 0.023 points for SSIM respectively. Based on these results, the FML verifies that aligning frequency components allow the HVDM to enhance temporal changes and perceptual quality of generated videos.
HVDM (w/o FML) 8.08 0.041 33.67 0.892
HVDM (w/0 FML) 5.35 0.038 34.00 0.915
Table 5: Ablation study on frequency matching loss.
@@ 6 Conclusion
We proposed a novel hybrid video autoencoder for video generation using diffusion model, dubbed HVDM. We investigated the effectiveness of our hybrid architecture that combines 2D projected and 3D volume representation, leveraging the strengths of frequency matching loss. By incorporating these representations with spatio-temporal cross-attention, HVDM not only address the challenges inherent in video generation, such as the high dimensionality and complexity of video data but also generates high-quality videos with improved realism. Our extensive experiments on benchmark datasets demonstrated the superior performance, compared to the other baseline methods. Furthermore, our approach allows for a wide range of video applications, such as long video generation, image-to-video, and video dynamics control, showing the flexibility and robustness of our method.
Appendix 0.A Implementation Details
In this section we describe details on the model configurations and hyper-parameters used for training of our HVDM. Specifically, for 2D triplane and 3D wavelet representation of our hybrid autoencoder, we modify existing PVDM [53] autoencoder architecture: 1) we use reduced depth of each layer in the transformer, 2) we use the decomposed wavelet input instead of the raw video, 3) we use 3D convolution networks to encode the decomposed wavelet features, and 4) we add cross-attention module to fuse 2D projected and 3D wavelet latent. As for the diffusion model, we use a simple 3D convolutional U-Net architecture [10]. We provide the detailed information for our hybrid autoencoder in Tab. 6 and U-Net in Tab. 7.
UCF-101 [38] SkyTimelapse [51] TaiChi [33]
@@ Appendix 0.A Implementation Details
Table 6: Autoencoder configuration and training details.
UCF-101 [38] SkyTimelapse [51] TaiChi [33]
@@
Table 7: Our denoising 3D U-Net configuration and training details.
Appendix 0.B Experiment Setup
0.B.1 Datasets
To provide a comprehensive evaluation with baseline, we follow the settings of PVDM for UCF-101 [38] and SkyTimelapse [51] datasets and as well as the settings of LVDM [33] for the TaiChi [33] dataset. For training the model, we crop and resize all video clips to $256\times 256$ scale. All video clips are extracted as consecutive 16 or 128 frame lengths from a randomly selected video.
UCF-101 is an action recognition dataset comprising realistic action videos, including 101 action categories. Following the common practice for evaluating unconditional video generation on UCF-101 [38], we use the train split (9,357 videos) for training and test split (3,963 videos) for evaluation.
@@ 0.B.1 Datasets
TaiChi is a Tai-Chi action dataset containing 3,078 videos in total. For a fair comparison with the LVDM [14], we use the train+test splits for both model training and evaluation. Moreover, we reproduced PVDM [53] under the same setup as the model weight was unavailable.
0.B.2 Metrics
For fair comparison, we follows the evaluation protocol of StyleGAN-V [35]. Unlike previous evaluation protocol that preprocess video datasets with fixed-length clips, StyleGAN-V [35] propose to sample video data first and then randomly select fixed-length clips, particularly avoiding bias issues with extremely long videos. As for evaluation metrics, we measure the video generation quality with the following two metrics: Fréchet Video Distance (FVD) [42] and Kernel Video Distance (KVD) [41]. In addition, we report the reconstruction quality of the autoencoder with the following four metrics: Learned Perceptual Image Patch Similarity (LPIPS) [55], Peak Signal-to-Noise Ratio [21] (PSNR), SSIM (Structural Similarity Index) [19], and R-FVD which denote FVD between real and reconstructed samples.
0.B.3 Inference settings for evaluations
We assessed the performance of video autoencoders by evaluating the reconstruction quality, using identical samples from the SkyTimelapse [51], UCF-101 [38], and TaiChi [33] datasets. To compare the performance of our method with state-of-the-art diffusion-based video generative models, we fixed the number of DDIM [37] steps to a constant value of 100. Additionally we extract 2,048 video clips for computing real statistics and generate 2,048 samples for evaluating fake statistics.
Appendix 0.C Inference Time and Memory
@@
Inference (time/memory)
Length $\rightarrow$ 16 128
LVDM [14] 8.85/.9.0 N/A
PVDM [53] 9.16/.9.6 137.8/10.1
HVDM (Ours) 9.90/11.6 151.4/12.1
Table 8: Comparison of the inference time (Sec) and memory (GB) of various video diffusion models. N/A denotes that the values cannot be measured because the model is unavailable.
Appendix 0.D Limitations and Future Directions
Text-to-video generative models are quickly advancing, offering users exciting opportunities to unleash their creativity in new ways. Although we haven’t executed experiments on large-scale video dataset for text-to-video generation tasks due to resource limitations, we are believe that our framework is also effective in this task as well.
Recently, Lumiere [4] generates the entire temporal duration of the video at once achieving global temporal consistency. In contrast to existing video models which synthesize sub-sampled set of keyframes followed by temporal super-resolution, they train model to learn globally coherent motion by downsampling the video in multiple space-time scales. As a alternative to compressing long video frames without being limited by memory requirements, employing the multi-level wavelet transform as a pooling operation would be an interesting research direction. Furthermore, although HVDM demonstrates potential for efficient and effective video synthesis, further research avenues exist for its enhancement. For example, exploring the design of diffusion model architectures that are specialized for our hybrid latents or wavelet filter banks that provide a highly efficient signal decomposition in the video domain would be an intriguing further research direction.
Appendix 0.E Additional Results
0.E.1 Video Reconstruction
Figure 7: Reconstruction results of our hybrid autoencoder on the SkyTimelapse [51].
Figure 8: Reconstruction results of our hybrid autoencoder on the TaiChi [33].
Figure 9: Reconstruction results of our hybrid autoencoder on the UCF-101 [38].
0.E.2 Unconditional Video Generation